# Extracting the singularity of a logarithmic partition function to investigate quantum phase transitions

### v2

@@ Abstract
In principle, the logarithm of a partition function and its derivatives can be used to rigorously determine the nature of phase transitions. However, this approach is often ineffective in numerical calculations in practice, since resolving the singularity of a logarithmic partition function requires extremely high-precision data—especially when distinguishing weakly first-order phase transitions and deconfined quantum critical points. Based on the finite-size scaling hypothesis, we propose that the singular contribution to the logarithmic partition function in quantum phase transitions can be extracted by eliminating the analytic terms. This is achieved by computing the Rényi thermal entropy (RTE) and its derivative (DRTE). We have derived the scaling relations for RTE and DRTE, which successfully determine the phase transition points and critical exponents through data collapse in various quantum many-body systems. Notably, beyond their utility in locating quantum critical points, we find that DRTE is remarkably effective in detecting weakly first-order phase transitions, such as in JQJ-Q models, which were initially thought to be continuous deconfined quantum phase transitions. The discontinuity of these transitions is challenging to detect using traditional local physical quantities. This work provides a new and powerful paradigm for studying quantum many-body phase transitions.
@@
Introduction.— Quantum phase transitions, associated spontaneous symmetry breaking, and universality classes are central concepts in condensed matter and statistical physics [1, 2, 3]. The nature of a phase transition is determined by the singularity in the free energy or its derivatives: a first-order transition is marked by a discontinuity in the first derivative, while a continuous transition involves singular higher-order derivatives. In principle, an exact calculation of the partition function would allow one to identify the transition based on these singularities. However, in practice, detecting such features in numerical simulations is challenging, as it requires extremely high-precision data. This issue is particularly severe when distinguishing weakly first-order transitions from deconfined quantum critical points [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. A key difficulty arises because the singular contribution is often overwhelmed by a much larger, non-singular term that scales with the system volume. Extracting the singular part therefore remains a major challenge in computational physics.
@@
For a quantum phase transition (inverse temperature $\beta\rightarrow\infty$; in the realistic simulations, the finite-size system gap $\Delta_{L}\sim\frac{1}{L^{z}}$, $\beta\sim L^{z}$ is usually used for probing the ground state properties, in which $L$ is the system size and $z$ is the dynamical critical exponent [17] ), the singularity of the free energy can be equivalently studied through the logarithm of the partition function. According to the finite-size scaling hypothesis [18, 17], a singular observable $Q$ (not necessarily divergent) near the critical point $g_{c}$ scales with system size as:
@@
where $\Delta g=g-g_{c}$, $\nu$ is is the correlation length critical exponent, $\sigma$ is a scaling power and $\tilde{S}$ is a universal scaling function. Since the singular part of the free energy is scale-invariant ($\sigma=0$[19, 18], the logarithm of the partition function can be expressed as:
@@
where $Z$ is a partition function, $d$ is the spatial dimension, $L$ is the system length, inverse temperature $\beta\sim L^{z}$ and $a(g)\beta L^{d}$ represents the leading term obeying a volume law, and $\mathcal{A}_{sub}(g,L,\beta)=\beta(a_{1}(g)L^{d-1}+a_{2}(g)L^{d-2}+a_{3}(g)L^{d-3}+\cdots)$ accounts for sub-leading analytic contributions ($a_{i}$ can be zero or non-zero). The dominant volume-law term often obscures the singular contribution $\tilde{S}$, which explains the difficulty in detecting phase transitions numerically through such singularities.
To isolate the singular term $\tilde{S}$, one can construct a quantity in which the dominant analytic contributions cancel out. Inspired by the method used to extract topological entanglement entropy [20, 21], we conjecture that the difference between $\ln Z(2\beta)$ and $2\ln Z(\beta)$ ($\beta\sim L^{z}$) can serve this role. Specifically, we argue that Eq. (2) can be rewritten as $\ln Z(2\beta)=2a(g)\beta L^{d}+2\mathcal{A}_{sub}(g,L,\beta)+\tilde{S}$ and $2\ln Z(\beta)=2a(g)\beta L^{d}+2\mathcal{A}_{sub}(g,L,\beta)+2\tilde{S}$. Coincidentally, this difference corresponds exactly to the second-order Rényi thermal entropy (RTE), a fundamental quantity in information theory  [22, 23, 24, 25, 26, 27, 28]:
@@
In this work, using our newly developed bipartite reweight-annealing quantum Monte Carlo (QMC) method [29, 30, 31, 32, 33, 34, 35], we validate the conjecture that the Rényi thermal entropy (RTE) is a dimensionless quantity capturing only the singular contribution. Moreover, its derivative (DRTE) exhibits a more pronounced singularity near the quantum critical point (QCP). We propose universal scaling forms for both RTE and DRTE, which are confirmed through data collapse analyses at (2+1)-dimensional (D) QCPs in the O($N$) universality classes, with $1\leq N\leq 3$. Notably, because the leading non-singular term in $\ln Z$ is canceled in DRTE, it provides a powerful diagnostic for weakly first-order phase transitions, even in relatively small systems (up to $L=56$), which are significant challenges to resolve using conventional numerical approaches.
Models and method.— We investigate quantum spin models with continuous QCPs in the (2+1)D O($N$) universality classes, where $1\leq N\leq 3$. We also examine first-order transitions, with a focus on weakly first-order cases, including some debated deconfined quantum critical points. The model Hamiltonians are described in detail in the following sections.
The second-order RTE can be expressed in terms of the ratio of partition functions at two different temperatures: $S^{(2)}=-\ln\frac{Z(2\beta)}{Z(\beta)^{2}}$. We compute the partition functions using the bipartite reweight-annealing algorithm [29, 30] in combination with stochastic series expansion (SSE) QMC simulations [36, 17, 37, 38, 39, 40, 41]. Following a method analogous to that used for the derivative of Rényi entanglement entropy and negativity [29, 42], the derivative of the RTE with respect to a general parameter $J$ is given by:
$\frac{dS^{(n)}}{dJ}=\frac{1}{1-n}\bigg{[}-n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z(n\beta)}+n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z(\beta)}\bigg{]}$ (4)
Here, the two expectation values are evaluated separately using ensembles corresponding to $Z(n\beta)$ and $Z(\beta)$, respectively. While the RTE is calculated using the reweight-annealing QMC method, the derivative (DRTE) is directly obtained from the two expectation values in Eq. (4); see Supplementary Materials (SM) for details. Considering the dynamical critical exponents $z=1$ of the models studied in this work, the $\beta=L$ is set to approach the quantum critical regime.
Finite-size scaling form.—For a zero-temperature system with ground-state degeneracy $d_{e}$, we can derive that: $S^{(n)}=\frac{1}{1-n}\ln\frac{Z(n\beta)}{Z(\beta)^{n}}=\ln d_{e}$$\tilde{S}(\Delta gL^{1/\nu})$, where $\tilde{S}$ represents the critical corrections. As shown in Fig.1, $S^{(2)}$ effectively probes both the degeneracy of discrete symmetry-breaking phases and the tower of excited states (TOS) associated with continuous symmetry-breaking.
Near criticality, the logarithmic partition function consists of a dominant non-singular term $a(J)\beta L^{d}$ and a dimensionless singular term $\tilde{S}(\Delta gL^{1/\nu})$. The leading non-singular term obscures the critical behavior, making it difficult to extract the singular part directly from the partition function. More interestingly, as shown in Fig. 1, the curves of $S^{(2)}$ for different system sizes intersect at the critical point, indicates that the RTE is independent of system size, providing strong numerical evidence that the leading non-singular contributions in $2\ln Z(\beta)$ and $\ln Z(2\beta)$ are identical and cancel exactly in the definition of RTE. This cancellation isolates the dimensionless singular term. Moreover, the DRTE displays a more pronounced divergence near the QCPs, further highlighting its sensitivity to critical fluctuations.
According to the finite-size scaling hypothesis  [18, 17], any singular quantity near a critical point follows a scaling form as given in Eq.(1). In particular, the dimensionless nature of the RTE ($\sigma=0$), implies the scaling form:
@@
From Eq. (5), we can readily derive the scaling behavior of the DRTE:
$\partial S^{(n)}(\Delta g,L)/\partial\Delta g=\tilde{S}^{\prime}(\Delta gL^{1/\nu})L^{1/\nu}.$ (6)
@@
Figure 1: (a,c) Second-order RTE and its derivative DRTE, and (b,d) corresponding data collapse analysis near the QCP of the bilayer Ising–Heisenberg model [(2+1)D Ising universality]. We set $J=3.045$ and take $J_{z}$ as the tuning parameter. (e–h) Results for the dimerized anisotropic Heisenberg model with $\Delta=0.9$ [(2+1)D O(2) universality], using $J_{2}=2.1035$ and tuning $J_{1}$. (i–m) Results for the dimerized isotropic Heisenberg model with $\Delta=1$ [(2+1)D O(3) universality], using $J_{2}=1$ and tuning $J_{1}$. To achieve good data collapse, we include a subleading correction: $Q(g,L)=L^{k}[\tilde{S}((g-g_{c})L^{1/\nu})+bL^{-\omega}]$. $\omega$ is close to 1, and to reduce the numerical instability caused by increasing the number of fitting parameters, we set $\omega=1$. For $Q=S^{(2)}$, $k=0$. $k=1/\nu$ when $Q$ is DRTE.
In the following section, by expanding the scaling function $\tilde{S}(\Delta gL^{1/\nu})$ (or $\tilde{S}^{\prime}(\Delta gL^{1/\nu})$) into polynomials, we will prove our scaling formula near the (2+1)D O(1) ($Z_{2}$), O(2), and O(3) QCPs through perfect data collapsing in Fig.1.
Ising transition.— We study the bilayer Ising-Heisenberg model with a (2+1)D Ising QCP [43]. The Hamiltonian is given by:
@@
We set $J=3.045$ and use $J_{z}$ as the tuning parameter, studying the second-order RTE and DRTE near the QCP at $J_{z,c}=1$ for various lattice sizes. The results are shown in Figs. 1(a) and (c). The RTE curves for different system sizes intersect at the QCP, while the DRTE exhibits a clear peak at the critical point. We fit the scaling forms in Eqs. (5) and (6) to the RTE and DRTE data, respectively. The fits yield $J_{z,c}=0.9999(1)$, $\nu=0.62(1)$ from RTE, and $J_{z,c}=1.000(3)$, $\nu=0.63(4)$ from DRTE. These values are agreement with the known QCP $J_{z,c}=1$  [43] and the (2+1)D Ising universality class value $\nu=0.63012$  [44, 45, 43]. The resulting universal scaling functions $\tilde{S}(\Delta gL^{1/\nu})$ and $\tilde{S}^{\prime}(\Delta gL^{1/\nu})$, obtained from data collapse, are displayed in Figs. 1(b) and (d), respectively.
Continuous symmetry breaking transitions.— We then investigate QCPs of the dimerized AF Heisenberg model [46, 47, 48]. The Hamiltonian reads:
@@
in which $J_{2}>J_{1}>0$ and $D_{ij}=S_{i}^{x}S_{j}^{x}+S_{i}^{y}S_{j}^{y}+\Delta S_{i}^{z}S_{j}^{z}$ represents the nearest-neighbor AF Heisenberg interaction with anisotropy parameter $\Delta$. In the regime where $J_{2}/J_{1}\simeq 1$, the ground state exhibits AF order. For $0<\Delta<1$, the AF order parameter lies in the $xy$-plane breaking the O(2) symmetry. For $\Delta>1$, the order parameter aligns along the $z$-axis, breaking the $\mathbb{Z}_{2}$ spin flip symmetry. When $\Delta=1$, the interaction is isotropic, and the AF order breaks the O(3) symmetry. In the limit $J_{2}/J_{1}\gg 1$, the ground state is the gapped dimer phase. The QCP from the gapped phase to the AF order belongs to the (2+1)D O(2)/Ising/O(3) universality class when $0<\Delta<1$/ $\Delta>1$ / $\Delta=1$  [46, 47, 48].
@@
We evaluate the RTE and its derivative near the QCPs for both the easy-plane case ($\Delta=0.9$, O(2) criticality) and the isotropic case ($\Delta=1$, O(3) criticality), with results shown in Figs. 1(e,g) and (i,k), respectively. In each case, $J_{2}$ is fixed and $J_{1}$ is tuned across the transition, as detailed in the caption. Similar to the Ising case, the RTE curves for different system sizes intersect at the QCPs, and their derivatives exhibit clear peaks at these points.
Figure 2: We set $Q=1$ and use $J$ as the tuning parameter. (a) The derivative of the RTE: DRTE $\partial S^{(2)}/\partial J$. The inset shows data on the right side of the transition point. (b) and (c) show the data for the derivative of the partition function at temperatures $\beta$ and $2\beta$ near the transition of the checkerboard $J$-$Q$ model. (d–f) Data for the $J-Q_{3}$ model. (g–i) Data for the $J-Q_{2}$ model. The raw values of the partition function derivative vary significantly across system sizes (see SM), making direct comparison difficult. To resolve this, we normalize the data by $L^{3}$, corresponding to $\beta L^{2}=L^{3}$.
We apply the scaling forms in Eqs. (5) and (6) to data near the (2+1)D O(2) and O(3) QCPs. For the O(2) QCP, fitting the RTE yields $J_{1,c}=0.997(1)$, $\nu=0.67(2)$, while the DRTE gives $J_{1,c}=0.999(1)$, $\nu=0.66(2)$. These values are consistent with the known critical parameters $J_{1,c}=1$ and $\nu=0.6703$ [44, 48].

### v3

# Complete finite-size scaling theory of Rényi thermal entropy for second, first and weak first order quantum phase transitions

@@ Abstract
Establishing the nature of a quantum phase transition in finite-size simulations – whether continuous, first-order, or weak first-order – is a fundamental challenge in quantum many-body computation. Especially, the weak first-order phase transition is affected by a super large correlation length and always displays as a continuous critical point in simulated finite-sizes. The core difficulty lies in the fact that there is no effective finite-size theory to distinguish these phase transitions in the realistic simulations limited by the computational resource. In this work, we have fixed this problem by introducing a unified finite-size framework based on the Rényi thermal entropy (RTE) and its derivative (DRTE) to detect and characterize quantum phase transitions. We derive complete scaling theories for the RTE and DRTE at second-order, first-order, and weak first-order transitions, showing that the DRTE naturally isolates the singular part of the free energy and strengthens the characteristics of various phase transitions in finite sizes. Using quantum Monte Carlo simulations, we demonstrate accurate data collapse and extraction of critical exponents at (2+1)-dimensional O($N$) critical points. More importantly, the DRTE provides a smoking-gun signature of weak first-order transitions through a clear double-peak structure and a crossing at zero, which we unambiguously observe in debated deconfined quantum criticality candidates such as the JJQQ models. Our approach offers a general, unbiased, and numerically efficient tool for probing the universal properties of quantum phase transitions, resolving long-standing ambiguities between continuous and weak first-order scenarios.
I Introduction
@@
Quantum phase transitions, associated spontaneous symmetry breaking, and critical behaviors are central concepts in condensed matter and statistical physics [1, 2, 3]. The nature of a phase transition is determined by singularities in the free energy or its derivatives, as rigorously defined in standard textbooks: a first-order transition is characterized by a discontinuity in the first derivative of the free energy, while a continuous transition involves singular higher-order derivatives. In principle, an exact analytical calculation of the partition function would allow unambiguous identification of the transition type through these singularities. However, only a handful of systems admit exact solutions, and the vast majority must be studied numerically. In practice, detecting such singular features in numerical simulations is highly challenging, as it requires extremely high-precision data to resolve the behavior of derivatives—particularly in strongly correlated many-body systems.
A long-standing and fundamental challenge lies in identifying genuine deconfined quantum critical points (DQCP) beyond the Landau-Ginzburg-Wilson paradigm. This framework describes continuous transitions between incompatible symmetry-breaking phases, such as the antiferromagnetic (AF) Néel order—where spin symmetry is broken—and the valence-bond solid (VBS)—where lattice-rotation symmetry is broken [4, 5, 6]. Such continuous transitions exhibit unconventional critical behavior and may host emergent symmetries and fractionalized excitations [7, 8, 9, 10, 11, 12, 13, 14]. However, in concrete model realizations, it remains controversial whether candidate DQCPs represent true continuous transitions or are instead weak first-order. Notable examples include nonlinear sigma models with topological terms [15, 16, 17, 18], the JQJ-Q models [5, 19, 20, 21, 22, 23], and other bosonic or fermionic systems [24, 25, 26, 27].
The core difficulty stems from the fact that correlation lengths at these candidate DQCPs are often much longer than the system sizes accessible in simulations—even if the transition is weak first-order rather than genuinely critical. Consequently, the observed critical-like behavior closely mimics that of a continuous transition, making definitive discrimination extremely difficult. This phenomenon is understood within a physical picture in which the weak first-order transition occurs in close proximity to a critical point—though some researchers argue that the underlying criticality is described by a complex conformal field theory (CFT)[28, 29, 18, 30], while others propose it arises from a multicritical point[31, 32, 33]. As a result, in finite-size simulations, weak first-order transitions near a critical fixed point are nearly indistinguishable from continuous ones. Given current computational limits, reliably distinguishing between these scenarios across accessible system sizes remains a major challenge.
Crucially, there is currently no unified and effective finite-size scaling theory capable of robustly differentiating second-order and weak first-order transitions. As a consequence, most numerical studies lack a definitive “smoking-gun” signature for identifying the true nature of the transition. In this work, we address this gap by developing a complete scaling theory based on the Rényi thermal entropy (RTE) and its derivative (DRTE)—physical observables that naturally cancels the leading analytic contributions to the free energy, thereby isolating the singular part that governs critical behavior. Moreover, the DRTE exhibits distinct and universal signatures in both first-order and weak first-order transitions, enabling clear discrimination between different types of quantum phase transitions. In continuous and weak first-order cases, it allows for the accurate extraction of critical exponents through data collapse, where the extracted exponents in the latter reflect those of the nearest critical fixed point.
Thus, our framework enables the nature of quantum phase transitions to be unambiguously identified through finite-size scaling of the RTE and DRTE, providing a general, unbiased, and numerically efficient tool for resolving long-standing debates in quantum criticality.
II Second order phase transition
II.1 Finite-size scaling form
@@
Let us start from the second order phase transition. Since we discuss quantum phase transition, $\beta\sim L^{z}$ is required ($\beta=1/T$ is the inverse temperature, $L$ is the system size and $z$ is the dynamical critical exponent). The free energy $F=-\frac{1}{\beta}lnZ$ of a system can be replaced by a logarithm of the partition function $lnZ$ because $\beta$ is irrelevant here. As we know, due to the correlation length diverging at a second order phase transition point, the free energy $F$ or the logarithmic partition function $lnZ$ must have a singular term. According to the finite-size scaling hypothesis [34, 35], a singular observable $Q$ (not necessarily divergent) near the critical point $g_{c}$ scales with system size as:
@@
where $\Delta g=g-g_{c}$, $g$ is the tunable parameter of a quantum phase transition and $g_{c}$ is the parameter at the critical point. $\nu$ is the correlation length critical exponent, $\sigma$ is a scaling power and $\tilde{S}$ is a universal scaling function. Since the singular part of the free energy is scale-invariant ($\sigma=0$[36, 34], the logarithm of the partition function can be expressed as:
@@
where $Z$ is a partition function, $d$ is the spatial dimension, $L$ is the system length, inverse temperature $\beta\sim L^{z}$ and $a(g)\beta L^{d}$ represents the leading term obeying a volume law, and $\mathcal{A}_{sub}(g,L,\beta)=\beta(a_{1}(g)L^{d-1}+a_{2}(g)L^{d-2}+a_{3}(g)L^{d-3}+\cdots)$ accounts for sub-leading analytic contributions ($a_{i}$ can be zero). The dominant analytic terms often obscure the singular contribution $\tilde{S}$, which explains the difficulty in detecting phase transitions numerically through such singularities.
To isolate the singular term $\tilde{S}$, one can construct a quantity in which the dominant analytic contributions cancel out. Inspired by the method used to extract topological entanglement entropy [37, 38], we conjecture that the difference between $\ln Z(2\beta)$ and $2\ln Z(\beta)$ ($\beta\sim L^{z}$) can serve this role. Specifically, we argue that Eq. (2) can be rewritten as
$\ln Z(2\beta)=2a(g)\beta L^{d}+2\mathcal{A}_{sub}(g,L,\beta)+\tilde{S}$ (3)
and
$2\ln Z(\beta)=2a(g)\beta L^{d}+2\mathcal{A}_{sub}(g,L,\beta)+2\tilde{S}$, (4)
so that their difference $\ln Z(2\beta)-2\ln Z(\beta)]$ scales as $\sim\tilde{S}$. This is indeed physically reasonable, because for a quantum phase transition, temperature is an irrelevant field. As long as it is proportional to the finite-size energy gap ($\beta\sim L^{z}\rightarrow\infty$), different multiples should only affect the analytic terms.
We originally intended to assign a name to the difference between the two physical quantities. Coincidentally, this difference turns out to be exactly the second-order Rényi thermal entropy—also known as the purity—a fundamental quantity in information theory [39, 40, 41, 42, 43, 44, 45]:
@@
In particular, according to Eq.(1), the dimensionless nature of the RTE ($\sigma=0$), implies the scaling form:
@@
From Eq. (6), we can readily derive the scaling behavior of the derivative of RTE (DRTE):
$\frac{\partial S^{(n)}(\Delta g,L)}{\partial\Delta g}=\tilde{S}^{\prime}(\Delta gL^{1/\nu})L^{1/\nu}.$ (7)
@@
If our above analysis is correct, the RTE should have a crossing at a critical point since it is dimensionless like Binder ratio. And the DRTE can extract the critical exponent $\nu$ by data collapse. In the following section, by expanding the scaling function $\tilde{S}(\Delta gL^{1/\nu})$ (or $\tilde{S}^{\prime}(\Delta gL^{1/\nu})$) into polynomials: $\tilde{S}(x)=\sum_{k=0}^{k_{\mathrm{max}}}\frac{1}{k!}\tilde{S}_{0}^{(k+1)}(0)x^{k}$. we will numerically prove our scaling formula near the widely accepted (2+1)D O(1) ($Z_{2}$), O(2), and O(3) quantum critical points (QCPs) through perfect data collapse in Fig.1. The critical point $g_{c}$, the critical exponent $\nu$, and the coefficients $\tilde{S}_{0}^{(k+1)}(0)$ are the fitting parameters and $k_{\mathrm{max}}=6$ in the following numerical analysis.
Figure 1: (a,c) Second-order RTE and its derivative DRTE, and (b,d) corresponding data collapse analysis near the QCP of the bilayer Ising–Heisenberg model [(2+1)D Ising universality]. We set $J=3.045$ and take $J_{z}$ as the tuning parameter. (e–h) Results for the dimerized anisotropic Heisenberg model with $\Delta=0.9$ [(2+1)D O(2) universality], using $J_{2}=2.1035$ and tuning $J_{1}$. (i–m) Results for the dimerized isotropic Heisenberg model with $\Delta=1$ [(2+1)D O(3) universality], using $J_{2}=1$ and tuning $J_{1}$. To achieve good data collapse, we include a subleading correction: $Q(g,L)=L^{k}[\tilde{S}((g-g_{c})L^{1/\nu})+bL^{-\omega}]$. $\omega=0.830$, $0.789$, and $0.782$ correspond to the (2+1)D Ising, O(2), and O(3) critical points, respectively [46, 47, 48, 49]. For $Q=S^{(2)}$, $k=0$. $k=1/\nu$ when $Q$ is DRTE.
II.2 Numerical methods
The RTE can be expressed in terms of the ratio of partition functions at two different temperatures: $S^{(2)}=-\ln\frac{Z(2\beta)}{Z(\beta)^{2}}$. We compute the partition functions using the bipartite reweight-annealing algorithm [50, 51, 52] in combination with stochastic series expansion (SSE) QMC simulations [53, 35, 54, 55, 56, 57]. Following a method analogous to that used for the derivative of Rényi entanglement entropy and negativity [50, 58], the derivative of the RTE with respect to a general parameter $g$ of the Hamiltonian $H$ is given by:
$\frac{dS^{(n)}}{dg}=\frac{1}{1-n}\bigg[-n\beta\bigg\langle\frac{dH}{dg}\bigg\rangle_{Z(n\beta)}+n\beta\bigg\langle\frac{dH}{dg}\bigg\rangle_{Z(\beta)}\bigg]$ (8)
Here, the two expectation values are evaluated separately using ensembles corresponding to $Z(n\beta)$ and $Z(\beta)$, respectively. While the RTE is calculated using the reweight-annealing QMC method [59, 60, 61], the DRTE is directly obtained from the two expectation values in Eq. (8); see Appendix A for details. Considering the dynamical critical exponents $z=1$ of the models studied in this work, the $\beta=L$ is set to approach the quantum critical regime.
II.3 Ising transition
We study the bilayer Ising-Heisenberg model with a (2+1)D Ising QCP [62]. The Hamiltonian is given by:
@@ II.3 Ising transition
We set $J=3.045$ and use $J_{z}$ as the tuning parameter, studying the second-order RTE and DRTE near the QCP at $J_{z,c}=1$ for various lattice sizes. The results are shown in Figs. 1(a) and (c). The RTE curves for different system sizes intersect at the QCP, while the DRTE exhibits a clear peak at the critical point. We fit the scaling forms in Eqs. (6) and (7) to the RTE and DRTE data, respectively. The fits yield $J_{z,c}=0.9999(1)$, $\nu=0.629(7)$ from RTE, and $J_{z,c}=1.000(3)$, $\nu=0.63(4)$ from DRTE. These values are in agreement with the known QCP $J_{z,c}=1$  [62] and the (2+1)D Ising universality class value $\nu=0.63012$  [63, 64, 62]. The resulting universal scaling functions $\tilde{S}(\Delta gL^{1/\nu})$ and $\tilde{S}^{\prime}(\Delta gL^{1/\nu})$ with $\Delta g=J_{z}-J_{z,c}$, obtained from data collapse, are displayed in Figs. 1(b) and (d), respectively.
II.4 Continuous symmetry phase transitions
We then investigate QCPs of the dimerized AF Heisenberg model [65, 66, 67]. The Hamiltonian reads:
@@ II.4 Continuous symmetry phase transitions
in which $J_{2}>J_{1}>0$ and $D_{ij}=S_{i}^{x}S_{j}^{x}+S_{i}^{y}S_{j}^{y}+\Delta S_{i}^{z}S_{j}^{z}$ represents the nearest-neighbor AF Heisenberg interaction with anisotropy parameter $\Delta$. In the regime where $J_{2}/J_{1}\simeq 1$, the ground state exhibits AF order. For $0<\Delta<1$, the AF order parameter lies in the $xy$-plane breaking the O(2) symmetry. For $\Delta>1$, the order parameter aligns along the $z$-axis, breaking the $\mathbb{Z}_{2}$ spin flip symmetry. When $\Delta=1$, the interaction is isotropic, and the AF order breaks the O(3) symmetry. In the limit $J_{2}/J_{1}\gg 1$, the ground state is the gapped dimer phase. The QCP from the gapped phase to the AF order belongs to the (2+1)D O(2)/Ising/O(3) universality class when $0<\Delta<1$/ $\Delta>1$ / $\Delta=1$  [65, 66, 67]. For a zero-temperature system with ground-state degeneracy $d_{e}$, the RTE actually reflects the logarithmic degeneracy: $S^{(2)}=\ln d_{e}$. As shown in Fig.1, $S^{(2)}$ effectively probes both the degeneracy of discrete symmetry-breaking phases and the Anderson tower of states associated with continuous symmetry-breaking [68, 69, 12].
@@
We evaluate the RTE and its derivative near the QCPs for both the easy-plane case ($\Delta=0.9$, O(2) criticality) and the isotropic case ($\Delta=1$, O(3) criticality), with results shown in Figs. 1(e,g) and (i,k), respectively. In each case, $J_{2}$ is fixed and $J_{1}$ is tuned across the transition, as detailed in the caption. Similar to the Ising case, the RTE curves for different system sizes intersect at the QCPs, and their derivatives exhibit clear peaks at these points. We apply the scaling forms in Eqs. (6) and (7) to data near the (2+1)D O(2) and O(3) QCPs. For the O(2) QCP, fitting the RTE yields $J_{1,c}=0.997(1)$, $\nu=0.67(2)$, while the DRTE gives $J_{1,c}=0.9994(8)$, $\nu=0.66(1)$. These values are consistent with the known critical parameters $J_{1,c}=1$ and $\nu=0.6703$ [63, 67].